OpenStudy (anonymous):

using the concept of slope ,find t such that the three points are collinear: (-3,3) (t,-1) (8,6)

OpenStudy (anonymous):

collinear= same line, correct?

OpenStudy (anonymous):

Find the line, plug in -1 for x, and solve for y (which would be t)

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

slope= change in y over change in x

OpenStudy (anonymous):

3-6/-3-8

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

y=1/4 x+b

OpenStudy (anonymous):

Now plug in a point and solve for b

OpenStudy (anonymous):

3=1/4(-3)+b

OpenStudy (anonymous):

3=-3/4 +b

OpenStudy (anonymous):

3+3/4=b

OpenStudy (anonymous):

3 3/4 (aka 3.75)=b

OpenStudy (anonymous):

y=1/4 x +3.75

OpenStudy (anonymous):

Now plug in -1 for y and t for x

OpenStudy (anonymous):

-1=1/4(t)+3.75

OpenStudy (anonymous):

-4.75=1/4 t

OpenStudy (anonymous):

-4.75/(1/4)=t

OpenStudy (anonymous):

-19=t

OpenStudy (anonymous):

for the slope i got 3/11

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

6-3/8-(-3)

OpenStudy (anonymous):

Should be -6-3/8--3 -or- -3-8/3--6

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

second equation= 3--6/-3-8

OpenStudy (anonymous):

But either way we both goofed up :p slope is - 9/11

OpenStudy (anonymous):

But now that u know how to do the problem, follow the same steps only with - 9/11 as the slope

OpenStudy (anonymous):

ya i get the concept thanks for your help